OpenStudy (anonymous):

Two objects are moving along separate linear paths where each path is described by position, d, and time, t. The variable d is measured in meters, and the variable t is measured in seconds. The equation describing the graph of the position of the first object with respect to time is d = 2.5t + 2.2. The graph of the position of the second object is a parallel line passing through (t = 0, d = 1). What is the equation of the second graph?

OpenStudy (mathstudent55):

What do you know about the slopes of parallel lines?

OpenStudy (anonymous):

not very much i was out of school for a while in the hospital and i have to get this stuff dont to play football

OpenStudy (anonymous):

done"

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Slope is the tilt of a line. If lines are parallel, they have the same slope.

OpenStudy (anonymous):

Right

OpenStudy (mathstudent55):

In the equation d = 2.5t + 2.2, the 2.5 is the slope. Since the other path is parallel, it must have the same slope, 2.5. The other path must have an equation d = 2.5t + b where b is a number we will find.

OpenStudy (mathstudent55):

We are given a point of the second path, (0, 1) We replace t and d of the equation with t = 0 and with d = 1. d = 2.5t + b, and we solve for b. 1 = 2.5(0) + b 1 = 0 + b 1 = b We see that b = 1 Now we rewrite the equation of the second path using b = 1. d = 2.5t + 1 That is the answer.

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

You're welcome.